a stone of mass 0.2mg is thrown vertically up with a speed of 19.6m/s.calculate its potential and kinetic energy after 1 secbio_

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
The initial velocity is 19.6m/s so initial kinetic energy = 1/2(mv2)
=1/2(.2X10-3)(19.6)(19.6) =3.846mJ
Now after 1 sec velocity will be v= u-gt =19.6-9.8=9.8
so k.e now will be =1/2(.2X10-3)(9.8)(9.8)=.9604mJ 
since the enrgy is conserved to potential energy wil be =3.846-.9604=2.886mJ